OpenStudy (anonymous):

Find a cubic function with the given zeros. 6, -3, 5 A) f(x) = x3 - 8x2 + 3x + 90 B) f(x) = x3 - 8x2 - 3x + 90 C) f(x) = x3 + 8x2 - 3x + 90 D) f(x) = x3 - 8x2 - 3x - 90

OpenStudy (anonymous):

Becuase we have three zeros of the cubic function, we can say that : \[f(x)=(x-6)(x+3)(x-5)\] So, by developing the expression you will get the answer ;)

OpenStudy (anonymous):

I know that, But what do I do next?

OpenStudy (anonymous):

We have : \[\Large f(x)=(x-6)(x+3)(x-5)\\ \Large ~~=(x^2+3x-6x-18)(x-5)\\ \Large ~~=(x^2-3x-18)(x-5)\\ \Large~~=x^3-5x^2-3x^2+15x-18x+90\\ \Large~~=x^2-8x^2-3x+90 \]

OpenStudy (anonymous):

Thanks much!